# A Differentiable Alignment Framework for Sequence-to-Sequence Modeling via Optimal Transport

### v1

@@ 7 Conclusion and Future Work
Impact Statement
We introduce the OTTC framework, which improves alignment performance in speech processing by mitigating the peaky behavior of CTC models. Our method shows considerable gains in alignment accuracy across multiple datasets and holds great promise for applications in automatic transcription, speaker diarization, and medical speech analysis. In particular, it can aid clinicians working with ASR tools for patients with conditions like stuttering and dysarthria, where precise temporal alignment enables more reliable transcriptions. This enhanced alignment could help clinicians save time and resources by streamlining speech analysis, ultimately improving the efficiency and accessibility of ASR-based tools in the medical field.
@@ A.3 Supplementary Experimental Insights
A.3.1 Ablation Studies
@@
Comments on soft-DTW. In soft-DTW, only the first and last elements of sequences are guaranteed to align, while all in-between frames or targets may be ignored; i.e., there is no guarantee that soft-DTW will yield a discrete monotonic alignment. A “powerful” transformation $F$ can map $\mathbf{x}$ to $F($$\mathbf{x}$)))) in such a way that soft-DTW ignores the in-between transformed frames ($F($$\mathbf{x}$))))) and targets ($\mathbf{y}$), which we refer to as a collapse (Section 4.2.1). This is why transformations learned through sequence comparison are typically constrained (e.g., to geometric transformations like rotations) (Vayer et al., 2022). Since transformer architectures are powerful, they are susceptible to collapse as demonstrated by the following experiment we conducted using soft-DTW as the loss function. On the 360h-LibriSpeech setup with Wav2Vec2-large model, the best WER achieved using soft-DTW is 39.43%. In comparison, CTC yields 2.77% whereas the proposed OTTC yields 3.00%. A key advantage of our method is that, by construction, such a collapse is not possible.
A.3.2 Alignment Analysis
Temporal evolution of alignment. An example of the evolution of the alignment in the OTTC model during training for 40 epochs without freezing OT weights prediction head is shown in Figure 7. Note that during the initial phase of training, there is significant left/right movement of boundary frames for all groups. As training progresses, the movement typically stabilizes to around 1-2 frames. While this can be considered “relatively stable” in terms of alignment, the classification loss (i.e., cross-entropy) in the OTTC framework is still considerably affected by these changes. This change of the loss is what impacts the final performance and the performance difference between freezing or not-freezing the alignments.

### v2

@@ A.3 Supplementary Experimental Insights
A.4 Note on Dynamic Time Warping (DTW)
It is important to highlight the distinction between our approach and DTW-based [48] alignment methods, particularly the differentiable variations such as soft-DTW [55]. These methods generally have quadratic complexity [55], making them significantly more computationally expensive than ours. Furthermore, in DTW-based methods, the alignment emerges as a consequence of the sequences themselves. When the function $F$ is powerful, the model can collapse by generating a sequence $F(\{{\bm{x}}\}_{n})$ that induces a trivial alignment [56] (see Appendix A.4.1, where we conducted experiments using soft-DTW for ASR to illustrate this). To mitigate this issue, regularization losses [56, 57] or constraints on the capacity of $F$ [58, 59] are commonly introduced. However, using regularization losses lacks theoretical guarantees and introduces additional hyperparameters. Furthermore, constraining the capacity of $F$, although more theoretically sound, makes tasks requiring powerful encoders on large datasets impractical. In contrast, our method decouples the computation of the alignment from the transformation function $F$, offering more flexibility to the model as well as built-in temporal alignment constraints and theoretical guarantees against collapse.
A.4.1 Ablation Studies
@@
Comments on soft-DTW. In soft-DTW, only the first and last elements of sequences are guaranteed to align, while all in-between frames or targets may be ignored; i.e., there is no guarantee that soft-DTW will yield a discrete monotonic alignment. A “powerful" transformation $F$ can map $\mathbf{x}$ to $F($$\mathbf{x}$)) in such a way that soft-DTW ignores the in-between transformed frames ($F($$\mathbf{x}$))) and targets ($\mathbf{y}$), which we refer to as a collapse (Section 4.2.1). This is why transformations learned through sequence comparison are typically constrained (e.g., to geometric transformations like rotations) [58]. Since transformer architectures are powerful, they are susceptible to collapse as demonstrated by the following experiment we conducted using soft-DTW as the loss function. On the 360h-LibriSpeech setup with Wav2Vec2-large model, the best WER achieved using soft-DTW is 39.43%. In comparison, CTC yields 2.77% whereas the proposed OTTC yields 3.00%. A key advantage of our method is that, by construction, such a collapse is not possible.
A.4.2 Alignment Analysis
Temporal evolution of alignment.
An example of the evolution of the alignment in the OTTC model during training for 40 epochs without freezing OT weights prediction head is shown in Figure 7. Note that during the initial phase of training, there is significant left/right movement of boundary frames for all groups. As training progresses, the movement typically stabilizes to around 1-2 frames. While this can be considered “relatively stable" in terms of alignment, the classification loss (i.e., cross-entropy) in the OTTC framework is still considerably affected by these changes. This change of the loss is what impacts the final performance and the performance difference between freezing or not-freezing the alignments.
@@ A.4.2 Alignment Analysis
Appendix B Limitations
The primary limitation of the current work is the observed trade-off between significantly improved alignment quality and a higher WER in ASR tasks compared to CTC. Further research is necessary to bridge this transcription accuracy gap. Additionally, the framework’s performance, particularly the quality of learned alignments and ASR accuracy, can be sensitive to the configuration of label weights ($\bm{\beta}_{q}$). The current use of fixed, uniform weights is a simplification, and developing strategies to learn $\bm{\beta}_{q}$ or devise more adaptive approaches without encountering degenerate solutions or overly complex training dynamics remains an area for future exploration. Finally, while the SOTD framework and OTTC loss show promise, their empirical validation and necessary adaptations have been primarily focused on ASR, with extensive investigation for a broader range of sequence-to-sequence tasks still required.
Appendix C Broader Impacts
Our work has the potential to positively impact several application areas. Improved temporal alignment can benefit domains such as medical speech analysis (e.g., detecting pathological cues), language learning tools (e.g., pronunciation feedback), and real-time captioning systems (e.g., enhanced synchronization for accessibility). The proposed methodology also advances sequence modeling by introducing a more interpretable alignment mechanism.
However, responsible deployment remains essential. The current trade-offs in transcription accuracy must be carefully considered before applying this approach in high-stakes scenarios. Additionally, as with all ASR technologies, there is a risk of biased performance across different demographic groups or speaking styles. Future work should address these concerns by incorporating fairness and robustness considerations. The interpretability gained from a single, learned alignment path may also support transparency and error analysis.